# AI Agents in Financial Markets: Architecture, Applications, and Systemic Implications

### 14 Mar 2026 v1

@@ Abstract
Recent advances in large language models, tool-using agents, and financial machine learning are transforming financial automation from isolated prediction tasks into integrated decision systems that can perceive information, reason over objectives, and generate or execute actions. As AI capabilities expand, financial institutions are beginning to deploy such systems across trading, portfolio management, risk monitoring, and decentralised finance. Yet the emerging literature remains fragmented across domain-specific language models, agent architectures, trading applications, and regulatory analyses, leaving the broader market implications of AI agents insufficiently understood.
This paper develops an integrative framework for analysing agentic finance: financial market environments in which autonomous or semi-autonomous AI systems participate in information processing, decision support, and execution workflows. The analysis proceeds in three steps. First, the paper proposes a four-layer architecture of financial AI agents covering data perception, reasoning engines, strategy generation, and execution with control. Second, it introduces the Agentic Financial Market Model (AFMM), which links agent design parameters-such as autonomy depth, model heterogeneity, execution coupling, infrastructure concentration, and supervisory observability-to market-level outcomes including efficiency, liquidity resilience, volatility, and systemic risk. Third, the paper outlines and illustrates an empirical research agenda for evaluating agentic systems using publicly observable financial and textual data.
The central argument is that the systemic implications of AI in finance depend less on model intelligence alone than on how agent architectures are distributed, coupled, and governed across financial institutions. In the near term, the most plausible equilibrium is likely to involve bounded autonomy, where AI agents operate as supervised co-pilots, monitoring systems, and constrained execution modules embedded within human decision processes. Understanding how these systems reshape financial market dynamics therefore represents an important agenda for both financial research and regulatory policy.
@@ 3. RQ3:
Methodologically, the paper is conceptual but explicitly research-oriented. It is not intended as a purely descriptive survey. Rather, it provides a theory-building framework that can guide future empirical work. The later sections therefore also outline an illustrative empirical application showing how the proposed framework can be operationalised using observable financial and textual data.
@@ 6.4 DeFi and on-chain intelligence
Table 1: Illustrative application map for AI agents in finance.
@@ 8.3 Variables and model specification
where $f_{ik}$ captures the frequency or salience of keyword $k$ in firm $i$’s 10-K or 10-Q filings. In the first-pass implementation, the keyword set includes terms related to the legacy core stack (COBOL, mainframe, z/OS, DB2, CICS), modernization processes (application modernization, cloud migration, replatforming, technical debt), and financial-system context (core banking, payment processing, policy administration). The baseline score uses simple term weights; a richer version could replace this with inverse-document-frequency or contextual NLP weights.
@@
The design considers narrow and moderate event windows, including $[0,+1]$, [1,+1][-1,+1], and [3,+3][-3,+3], with an estimation window of approximately [120,20][-120,-20] trading days. Because the February 24, 2026 event generates the strongest market response in the data, the baseline cross-sectional specification focuses on that event:
@@
where $Vendor_{i}$ identifies legacy-service providers and $Financial_{i}$ identifies financial-sector firms. This is intentionally a reduced-form specification. The objective is not to claim a complete structural model, but to test whether the sign and magnitude of repricing differ systematically by exposure and business model.
Construct Illustrative proxy Public or low-cost source
@@
Table 2: Illustrative data structure for the empirical application.
@@ 8.5 Illustrative findings
To move beyond design and provide a first empirical read, the application was implemented on a 31-firm sample using publicly retrievable price and filing data between June 2025 and March 2026. The sample comprises eight legacy-service vendors, sixteen legacy-heavy financial users, and seven software-or-infrastructure control firms. One candidate firm from the initial seed universe was dropped because the public market-data endpoint did not provide a stable history for the relevant window.
Table 3 reports subgroup means for the three focal dates. The clearest pattern appears around the February 24, 2026 market-validation event. For legacy-service vendors, mean CAR[1,+1]CAR[-1,+1] is approximately $6.39\%$. For legacy-heavy financial users, the corresponding figure is $2.61\%$, while the control portfolio is approximately zero by construction because it is also the benchmark. This is the central empirical result of the section: the market reaction is not uniform, and the sharpest negative repricing is concentrated in the group whose revenues are most directly threatened by accelerated modernization.
@@
E1: service vendors 8 0.0222-0.0222 0.0116-0.0116
E1: financial users 16 0.0097-0.0097 $0.0952$
E1: controls 7 $0.0000$ $0.1733$
E3: service vendors 8 0.0639-0.0639 $0.7296$
E3: financial users 16 0.0261-0.0261 $0.3668$
E3: controls 7 $0.0000$ $0.6966$
E4: service vendors 8 0.0004-0.0004 $0.3804$
E4: financial users 16 0.0306-0.0306 $0.2311$
E4: controls 7 $0.0000$ $0.5154$
Table 3: Illustrative event-study results for the AI-agent capability-shock application. Event E1 is the January 27, 2026 Anthropic service-delivery signal, E3 is the February 24, 2026 market-validation event, and E4 is the March 5, 2026 labour-market report. The February 23 media-amplification date is treated as overlapping with E3 and is therefore omitted from the headline table.
Table 4 sharpens this interpretation. In the February 24 cross section, the coefficient on the legacy-service vendor indicator is 0.0262-0.0262 with a $t$-statistic of approximately 1.95-1.95, suggesting economically meaningful underperformance for firms whose business models are more exposed to legacy maintenance and modernization revenues. By contrast, the filing-based exposure score is negative but not precisely estimated in this small sample, while the generic financial-sector dummy is close to zero. The empirical implication is important: the market does not appear to price all legacy exposure in the same way. What matters is not simply whether a firm mentions legacy technology, but whether its revenue model is expected to lose or gain from agentic modernization.
Regressor Coefficient $t$-statistic
Legacy exposure score 0.0005-0.0005 0.95-0.95
Legacy-service vendor dummy 0.0262-0.0262 1.95-1.95
Financial-sector dummy $0.0007$ $0.06$
@@
Table 4: Reduced-form cross-sectional regression for CAR[1,+1]CAR[-1,+1] around the February 24, 2026 market-validation event.
Taken together, these findings are consistent with the core claim of the AFMM. Capability disclosures about agentic AI appear to trigger market-wide attention, but price effects are heterogeneous and tied to the expected distribution of rents across exposed actors. In that sense, the application supports the broader thesis of the paper: the economic significance of agentic AI depends not only on abstract model capability, but on how specific architectures and workflows map into incumbent business models, modernization costs, and institutional dependencies.
@@ 8.6 Why the application matters for the full paper
In AFMM terms, the application captures a change in perceived autonomy depth ($A$), workflow substitutability, and execution-relevant capability before large-scale institutional diffusion is directly measurable. It therefore functions as a bridge between the architecture discussion in Section 4 and the systemic implications developed in Section 5. If capability news can already reprice adjacent firms through expectations alone, then widespread deployment could plausibly reshape market structure, business-model risk, and concentration dynamics at a larger scale.
@@ 9.1 From capability shocks to market structure
The paper’s main claim is not that a single announcement can reveal the full long-run impact of agentic finance. The point is narrower and more defensible. If secondary markets already differentiate across firms when new agentic capabilities become credible, then architecture and workflow design are economically relevant before full deployment is observable. The application therefore complements the conceptual framework by showing one concrete transmission channel from agent design to market outcomes: public capability news changes expected rents, which in turn changes valuation.
@@ 9.3 Limitations and next steps
Despite these limitations, the framework now has a clearer empirical anchor. It offers a plausible way to connect architecture, adoption, governance, and repricing within a single research agenda for agentic finance.
@@ 10 Conclusion
The empirical application reinforces this broader argument. A first-pass event-study design around AI-agent capability disclosures suggests that market attention rises broadly, but repricing is heterogeneous and concentrated in business models that appear more exposed to workflow substitution and modernization pressure. In that sense, the empirical illustration supports the paper’s central claim: the financial significance of agentic AI depends on how capabilities interact with institutional roles, infrastructure dependencies, and the distribution of economic rents.

### 14 Apr 2026 v2

@@ Abstract
Recent advances in large language models, tool-using agents, and financial machine learning are shifting financial automation from isolated prediction tasks to integrated decision systems that can perceive information, reason over objectives, and generate or execute actions. This paper develops an integrative framework for analysing agentic finance: financial market environments in which autonomous or semi-autonomous AI systems participate in information processing, decision support, monitoring, and execution workflows. The analysis proceeds in three steps. First, the paper proposes a four-layer architecture of financial AI agents covering data perception, reasoning engines, strategy generation, and execution with control. Second, it introduces the Agentic Financial Market Model (AFMM), a stylised agent-based representation linking agent design parameters such as autonomy depth, heterogeneity, execution coupling, infrastructure concentration, and supervisory observability to market-level outcomes including efficiency, liquidity resilience, volatility, and systemic risk. Third, it presents an illustrative empirical application based on event studies of AI-agent capability disclosures and heterogeneous market repricing. The paper argues that the systemic implications of AI in finance depend less on model intelligence alone than on how agent architectures are distributed, coupled, and governed across institutions. The empirical application is intentionally exploratory: it does not validate the full AFMM, but shows how one observable expectations channel can be studied using public data. In the near term, the most plausible equilibrium is bounded autonomy, in which AI agents operate as supervised co-pilots, monitoring systems, and constrained execution modules embedded within human decision processes.
@@ 3. RQ3:
Methodologically, the paper is conceptual but explicitly research-oriented. It is not intended as a purely descriptive survey. Rather, it provides a theory-building framework that can guide future empirical work. The later sections therefore also develop an illustrative empirical application showing how a limited but observable part of the framework can be operationalised using public financial and textual data.
@@ 2.1 Six streams of research on agentic finance
Recent review articles also confirm how quickly this space is maturing. New surveys synthesize the rise of agentic AI architectures in general and AI-agent applications in finance and fintech more specifically (Abou Ali et al., 2026; Rizinski and Trajanov, 2025; Aldridge et al., 2025). Those papers are useful reference points for the field. However, their primary objective is synthesis: they review architectures, use cases, and regulatory themes across a rapidly expanding domain. The present paper instead aims to provide a finance-specific conceptual framework that links agent architecture, institutional deployment, governance constraints, and market-level consequences, while also translating those design choices into a stylised market model.
@@ 2.3 Theoretical contribution of this paper
This positioning also helps distinguish the paper from adjacent literatures more sharply. Unlike review articles on AI or LLMs in finance, the paper does not primarily synthesise task-level findings. Unlike finance-specific LLM studies, it is not centred on benchmark performance in summarisation, question answering, or prediction. Unlike general agent-architecture papers, it focuses on finance-specific delegation boundaries, market structure, and governance constraints. Unlike the classical agent-based finance literature, it does not begin from stylised behavioural trading rules alone, but from modern AI-enabled decision pipelines embedded in institutions, data infrastructures, and supervisory regimes. The paper’s intended contribution is therefore conceptual integration: to connect agent design, institutional deployment, and market-level consequences within a single finance-oriented analytical framework.
@@ 6.4 DeFi and on-chain intelligence
Table 1: Application map for AI agents in finance.
@@ 7.2 Policy responses and supervisory strategies
In Europe, several recent regulations provide partial legal anchors for this agenda even though none was designed specifically for agentic finance. The Digital Operational Resilience Act (DORA) strengthens requirements around operational resilience, ICT risk management, and oversight of critical third-party technology providers (European Union, 2022). The Markets in Crypto-Assets Regulation (MiCA) shows how digital-asset infrastructures are increasingly being brought within explicit governance, disclosure, and conduct frameworks (European Union, 2023). The EU AI Act extends this trajectory by imposing risk-based obligations around high-impact AI systems, documentation, monitoring, and human oversight (European Union, 2024). Recent practitioner-oriented work on agentic investment firms points in a similar direction by stressing that architectural delegation, compliance automation, and supervisory integration must be designed together rather than treated as separate afterthoughts (Joshi, 2026). Taken together, these frameworks do not yet solve the governance problem of agentic finance, but they illustrate the direction of travel: resilience, traceability, outsourcing dependence, and human accountability are moving from soft principles toward harder institutional requirements.
@@ 8.2 Research question, sample construction, and exposure taxonomy
The implementation begins from a 32-name seed universe and retains 31 firms in the final event-study sample. The three retained groups are eight legacy-service vendors, sixteen legacy-heavy financial users, and seven control software or infrastructure firms. Fiserv (FI) was part of the initial seed universe but was excluded from the final event-study sample because the public daily-price endpoint used in the reproducibility pack did not provide a sufficiently stable history over the relevant estimation and event windows. This distinction between the seed universe and the final estimation sample matters for replication and is reported explicitly because the reduced-form cross-sectional results are based on the 31-firm sample rather than the larger candidate list.
Sample layer NN Firms in the final event-study sample
Legacy-service vendors 8 IBM, KD, DXC, ACN, CTSH, EPAM, GIB, ORCL
Legacy-heavy financial users 16 JPM, BAC, C, WFC, USB, PNC, AXP, COF, AIG, MET, PRU, TRV, FIS, GPN, ACIW, JKHY
Control software/infrastructure firms 7 MSFT, AMZN, NOW, CRM, ADBE, INTU, SNOW
Table 2: The 31-firm event-study sample used in the empirical application. The initial seed universe also contained Fiserv (FI), which was excluded because the public market-data endpoint used for the reproducibility pack did not provide a stable history for the relevant estimation and event windows.
@@ 8.3 Variables and model specification
where $f_{ik}$ captures the frequency or salience of keyword $k$ in firm $i$’s 10-K or 10-Q filings. In the implementation reported here, the dictionary is divided into three weighted blocks: legacy core-stack terms such as COBOL, mainframe, z/OS, DB2, and CICS; modernization and transformation terms such as application modernization, cloud migration, replatforming, legacy system(s), and technical debt; and financial-system context terms such as core banking, payment processing, policy administration, and claims processing. The baseline score uses prespecified term weights drawn from this dictionary rather than machine-learned weights. A richer version could replace this with inverse-document-frequency or contextual NLP weights, but the present design intentionally prioritises transparency over optimisation.
@@
The design considers narrow and moderate event windows, including $[0,+1]$, [1,+1][-1,+1], and [3,+3][-3,+3], with an estimation window of approximately [120,20][-120,-20] trading days. Because the February 24, 2026 market-validation event is the cleanest dissemination date and generates the strongest cross-sectional separation in the data, the baseline cross-sectional specification focuses on that event:
@@
where $Vendor_{i}$ identifies legacy-service providers and $Financial_{i}$ identifies financial-sector firms. The regression is estimated by ordinary least squares and reported with conventional standard errors. This is intentionally a reduced-form specification. The objective is not to claim a complete structural model, but to test whether the sign and magnitude of repricing differ systematically by exposure and business model.
Construct Baseline proxy Public or low-cost source
@@
Table 3: Data structure used in the empirical application.
Dictionary block Representative terms retained in the baseline score Typical weight range
Legacy core stack COBOL; mainframe; z/OS; DB2; CICS; legacy code 3.0–4.0
Modernization and transformation application modernization; cloud migration; replatforming; legacy system(s); technical debt; application migration 2.0–3.0
Financial-system context core banking; payment processing; policy administration; claims processing 1.5–2.5
Table 4: Keyword dictionary used to construct the legacy-modernization exposure score. The machine-readable reproducibility pack contains the full keyword-weight mapping.
@@ 8.5 Illustrative findings
To move beyond design and provide a first empirical read, the application was implemented on a 31-firm sample using publicly retrievable price and filing data between June 2025 and March 2026. The sample comprises eight legacy-service vendors, sixteen legacy-heavy financial users, and seven software-or-infrastructure control firms. As noted above, Fiserv (FI) was part of the 32-name seed universe but was excluded from the final event-study implementation because the public endpoint used for daily prices did not provide a sufficiently stable history for the relevant estimation window.
Table 5 reports subgroup means for the three focal dates together with standard errors in parentheses. The clearest pattern appears around the February 24, 2026 market-validation event. For legacy-service vendors, mean CAR[1,+1]CAR[-1,+1] is approximately $6.39\%$ (s.e. $0.0140$). For legacy-heavy financial users, the corresponding figure is $2.61\%$ (s.e. $0.0057$), while the control portfolio is approximately zero by construction because it is also the benchmark. The result should be read as an exploratory empirical illustration rather than a decisive validation exercise: the market reaction is not uniform, and the sharpest negative repricing is concentrated in the group whose revenues are most directly threatened by accelerated modernization.
@@
E1: service vendors 8 0.0222-0.0222 $(0.0054)$ 0.0116-0.0116
E1: financial users 16 0.0097-0.0097 $(0.0056)$ $0.0952$
E1: controls 7 $0.0000$ $(0.0102)$ $0.1733$
E3: service vendors 8 0.0639-0.0639 $(0.0140)$ $0.7296$
E3: financial users 16 0.0261-0.0261 $(0.0057)$ $0.3668$
E3: controls 7 $0.0000$ $(0.0058)$ $0.6966$
E4: service vendors 8 0.0004-0.0004 $(0.0072)$ $0.3804$
E4: financial users 16 0.0306-0.0306 $(0.0048)$ $0.2311$
E4: controls 7 $0.0000$ $(0.0119)$ $0.5154$
Table 5: Event-study results for the AI-agent capability-shock application. Event E1 is the January 27, 2026 Anthropic service-delivery signal, E3 is the February 24, 2026 market-validation event, and E4 is the March 5, 2026 labour-market report. Standard errors for subgroup mean CAR[1,+1]CAR[-1,+1] are reported in parentheses. The February 23 media-amplification date is treated as overlapping with E3 and is therefore omitted from the main table.
Table 6 sharpens this interpretation. In the February 24 cross section, the coefficient on the legacy-service vendor indicator is 0.0262-0.0262 with a standard error of $0.0134$ and a $t$-statistic of approximately 1.95-1.95, suggesting economically meaningful underperformance for firms whose business models are more exposed to legacy maintenance and modernization revenues. By contrast, the filing-based exposure score is negative but not precisely estimated in this small sample, while the generic financial-sector dummy is close to zero. The empirical implication is narrower than a full test of the AFMM: the market does not appear to price all legacy exposure in the same way. What matters in this exploratory application is not simply whether a firm mentions legacy technology, but whether its revenue model is expected to lose or gain from agentic modernization.
Regressor Coefficient Standard error $t$-statistic
Legacy exposure score 0.0005-0.0005 $0.0005$ 0.95-0.95
Legacy-service vendor dummy 0.0262-0.0262 $0.0134$ 1.95-1.95
Financial-sector dummy $0.0007$ $0.0130$ $0.06$
@@
Table 6: Cross-sectional regression for CAR[1,+1]CAR[-1,+1] around the February 24, 2026 market-validation event. The regression is estimated by ordinary least squares using the 31-firm implementation sample.
Table 7 adds a minimal robustness check by showing how subgroup means vary across alternative event windows for the February 24 market-validation event. The broad ordering is stable: legacy-service vendors remain the most negatively repriced subgroup, legacy-heavy financial users are less negative, and the control portfolio stays close to zero by construction. This is still a narrow robustness exercise, not a substitute for a larger multi-event design, but it helps show that the central pattern is not an artifact of a single event-window choice.
Subgroup $CAR[0,+1]$ CAR[1,+1]CAR[-1,+1] CAR[3,+3]CAR[-3,+3]
Legacy-service vendors 0.0092-0.0092 0.0639-0.0639 0.0763-0.0763
Legacy-heavy financial users $0.0026$ 0.0261-0.0261 0.0480-0.0480
Controls $0.0000$ $0.0000$ $0.0000$
Table 7: Subgroup mean cumulative abnormal returns across alternative event windows around the February 24, 2026 market-validation event.
Taken together, these findings are consistent with one observable implication of the AFMM rather than with a full empirical confirmation of the framework. Capability disclosures about agentic AI appear to trigger market-wide attention, but price effects are heterogeneous and tied to the expected distribution of rents across exposed actors. In that sense, the application is best read as an exploratory empirical illustration of the broader thesis of the paper: the economic significance of agentic AI depends not only on abstract model capability, but on how specific architectures and decision pipelines map into incumbent business models, modernization costs, and institutional dependencies.
@@ 8.6 Why the application matters for the full paper
In AFMM terms, the application captures a change in perceived autonomy depth ($A$), workflow substitutability, and execution-relevant capability before large-scale institutional diffusion is directly measurable. It therefore functions as a bridge between the architecture discussion in Section 4 and the systemic implications developed in Section 5. If capability news can already reprice adjacent firms through expectations alone, then widespread deployment could plausibly reshape market structure, business-model risk, and concentration dynamics at a larger scale. At the same time, this remains a single-channel exercise: it does not test the full AFMM, identify long-run equilibrium effects, or recover the internal decision logic of deployed institutional agents.
@@ 9.1 From capability shocks to market structure
The paper’s main claim is not that a single announcement can reveal the full long-run impact of agentic finance. The point is narrower and more defensible. If secondary markets already differentiate across firms when new agentic capabilities become credible, then architecture and decision-pipeline design are economically relevant before full deployment is observable. The application therefore complements the conceptual framework by showing one concrete transmission channel from agent design to market outcomes: public capability news changes expected rents, which in turn changes valuation.
@@ 9.3 Limitations and next steps
The current framework also abstracts from several issues that deserve more direct treatment in future work. One is the possibility of behavioural biases or mode-specific failures in AI systems, including overreaction to salient narratives, brittle planning, or common failure modes across highly similar models. Another is the feedback loop between human decision-makers and AI-generated recommendations. In many practical settings, the relevant unit is not a stand-alone agent but a human–AI ensemble in which attention, escalation, and override patterns affect outcomes.
@@
Despite these limitations, the framework now has a clearer empirical anchor. It offers a plausible way to connect architecture, adoption, governance, and repricing within a single research agenda for agentic finance without claiming that one small event-study exercise exhausts the empirical agenda.
@@ 10 Conclusion
The empirical application reinforces this broader argument in a limited but useful way. A first-pass event-study design around AI-agent capability disclosures suggests that market attention rises broadly, but repricing is heterogeneous and concentrated in business models that appear more exposed to workflow substitution and modernization pressure. The empirical illustration is therefore consistent with, rather than dispositive proof of, the paper’s central claim that the financial significance of agentic AI depends on how capabilities interact with institutional roles, infrastructure dependencies, and the distribution of economic rents.
@@
Acknowledgments
During the preparation of this manuscript, the author used large language model tools to assist with literature organisation, drafting support, and LaTeX editing. All generated output was reviewed, revised, and validated by the author, who takes full responsibility for the final content.
Funding
This research received no external funding.
Institutional Review Board Statement
Not applicable.
Informed Consent Statement
Not applicable.
Data Availability Statement
Publicly available data were used in the illustrative empirical application. The sources discussed in the manuscript include SEC EDGAR filings, FRED macro-financial series, and publicly accessible market-data endpoints such as Yahoo Finance and Alpha Vantage. No new proprietary dataset was created for this study. The local code and derived files used for the illustrative application are available from the corresponding author upon reasonable request.
Conflicts of Interest
The author declares no conflicts of interest.